### v8

# Search for the production of dark gauge bosons in the framework of Einstein-Cartan portal in the simulation of proton-proton collisions at $\sqrt{s}=13.6$ TeV

Abstract
In the present work, we study the possible production of the heavy neutral dark gauge boson (A) candidates, which originated from a simplified model based on the Einstein-Cartan gravity, in association with dark matter. This study has been performed by studying events with dimuon plus missing transverse momentum produced in the simulated proton-proton collisions at the Large Hadron Collider, at 13.6 TeV center of mass energy and integrated luminosity of 52 fb-1 corresponding to the LHC run 3 circumstances during 2022 and 2023. We provide upper limits, in case no new physics has been discovered, on the masses of various particles in the model as (A), as well as the heavy mediator (torsion field).
preprint: APS/123-QED
I Introduction
The Standard Model of particle physics SM has been tested for more than 40 years [1], and its predictions agree very well with all experimental observations. However, the SM is nowadays considered as a low-energy manifestation of other theories realized at high energy, generically known as BSM (Beyond the Standard Model) theories [2]. One motivation for BSM physics is to have a unified theory for the electromagnetic, weak, and strong interactions, in a unique Grand Unified Theory (GUT) [29]. The Super-Symmetry (SUSY) attempts to also include gravitation lead to models with extra spatial dimensions. These BSM models typically predict the existence of new dark particles at the TeV scale and higher.
The existence of heavy neutral bosons (Z) is a feature of many extensions of the Standard Model. They arise in extended gauge theories, including grand unified theories (GUT) [30], and other models like left-right symmetric models (LRM) [31]. A specific case is the sequential standard model (SSM), in which the Z boson has the same coupling as the SM Z boson [32]. Model of extra dimensions like Randall and Sundrum model (RS) [7] predicts the existence of heavy Kaluza-Klein gravitons. Searches for these heavy dark neutral gauge bosons have been performed at the CMS and ATLAS experiments, at the Large Hadron Collider (LHC), with no evidence of their existence using the full RUN II period of the LHC data taking [33, 34].
[…]
Many searches for DM have been performed by analyzing the data collected by the CMS experiment during RUN II. These searches rely on the production of a visible object ”X”, which recoils against the large missing transverse momentum from the dark matter particles leaving a signature of ($\text{X}+E^{\text{miss}}_{T}$) in the detector [18]. The visible particle could be an SM particle like Z [19, 20], W bosons [21, 22] or jets [23, 24], photon [25, 26] or SM Higgs boson [27, 28].
The ATLAS and CMS collaborations have previously conducted searches for the massive extra neutral gauge boson Z decaying to dileptons. These searches are based on the Grand Unified Theory (GUT) and Supersymmetry [29, 30, 31, 32]. Despite extensive efforts using the LHC data from the full RUN II period, no evidence of the existence of Z has been found [33, 34]. Results from the CMS collaboration have excluded the existence of Z in the mass range between 0.6 and 5.15 TeV with 95% Confidence Level (CL), while the ATLAS experiment has excluded the mass range between 0.6 and 5.1 TeV.
The direct detection of dark matter in underground experiments [35, 36, 37, 38, 39] presents various constraints. The model suggests that the interaction between dark matter and electrons can effectively explore light-dark matter. Exclusion limits confidently eliminate the possibility of dark matter with masses exceeding 4 MeV [40].
Indirect detection searches for dark matter are currently very exciting. The annihilation of dark matter into leptons or hadrons in space can potentially produce a signal of thermal WIMPs. The most reliable constraints for GeV dark matter annihilation into various final states are provided by Fermi-Lat observations of dwarf spheroidal galaxies [41, 42], AMS-02 positron fluxes [43, 44], and CMB [45, 46, 47] limits. These results indicate that the WIMP dark matter mass window from 20 GeV up to the unitarity limit [48] of 100 TeV remains largely open [49, 50].
In this analysis, we present a search for dark neutral gauge bosons (A), which originated in a simplified model in the Einstein-Cartan portal, at the LHC simulated proton-proton collisions with 13.6 TeV center of mass energy corresponding to the LHC run 3 circumstances [52]. The topology of the studied simulated events is dimuon, from the decay of A, plus large missing transverse momentum attributed to dark matter. The signal of the above process is very similar to that coming from the decay of Charginos pair into the lepton pair plus MET, carried out by ATLAS [53].
In the following section II, the theoretical formalism of the $U(1)_{D}$ simplified model based on the Einstein-Cartan gravity and its free parameters are presented. Then the simulation techniques used for event generation for the signal and SM background samples are displayed in section III. Afterward, the selection cuts and the strategy of the analysis are explained in section IV. Finally, this analysis’s results and summary are discussed in sections V and VI respectively.
Figure 1: Feynman diagram for the simplified model based on Einstein-Cartan gravity; for the production of dark gauge boson (A) in association to dark matter ($\chi$) pair [17].
II The simplified model in the framework of Einstein-Cartan gravity
The analyzed simplified model is based on the Einstein-Cartan gravity, which has been discussed in [17], assumes the production of dark matter from proton-proton collisions at the LHC in addition to a new heavy neutral dark gauge boson A.
[…]
The interaction terms, in the effective Lagrangian, between the torsion field and Dirac fermion $(\psi)$, are given by [17]
$\bar{\psi}i\gamma^{\mu}(\partial_{\mu}+i\texttt{g}_{\eta}\gamma^{5}S_{\mu}+...)\psi.$ (1)
According to the previous equation, the torsion field interacts with both the Standard Model (SM) and hidden fermions through its axial-vector part, with a universal coupling of $\texttt{g}_{\eta}$ = 1/8 at the classical level. This indicates that despite the SM and dark sectors existing in separate realms, the torsion field’s universal coupling to any fermion is essential for generating dark-sector fermions. This occurs when SM fermions collide, leading to the subsequent radiation or decay of lighter dark-sector particles, producing dark gauge bosons with a significantly unsuppressed production cross-section. In the case of the dynamic torsion field, the classical value of $\texttt{g}_{\eta}$ = 1/8 becomes unstable due to quantum corrections, causing $\texttt{g}_{\eta}$ to vary with the energy scale [17]. Due to the torsion mass being in the TeV-scale regime, the $A^{\prime}$ is produced with a high boost, and MET from dark-sector fermions is large where the SM backgrounds are low [17]. For this reason we have chosen $\texttt{g}_{\eta}=0.2$.
[…]
$\bar{\chi}(i\gamma^{\mu}D_{\mu}-M_{\chi})\chi,$ (2)
where $D_{\mu}=\partial_{\mu}+i\texttt{g}_{\eta}\gamma^{5}S_{\mu}+i\texttt{g}_{D}A^{\prime}_{\mu}$, $M_{\chi}$ is the dark matter mass and $\texttt{g}_{D}$ is the coupling of dark gauge boson to dark matter. The value of the coupling constant $\texttt{g}_{D}$ is taken from [17], which is $\texttt{g}_{D}=1.2$.
Parameters Description optimized value
$\texttt{g}_{\eta}$ Coupling of torsion field to Dirac fermions       0.2 [17]
$\texttt{g}_{D}$ Coupling of dark gauge boson to dark matter       1.2 [17]
$M_{\chi}$ Mass of dark matter $(\chi)$       $500$ GeV [50, 51]
$M_{A^{\prime}}$ Mass of dark gauge boson $(A^{\prime})$       $[200,...,900]$ GeV
$M_{ST}$ Mass of torsion field $(ST)$       $[1250,...,7000]$ GeV
Mass assumption $M_{A^{\prime}}<2M_{\chi}$ [17]
Table 1: The simplified model parameters based on Einstein Cartan gravity and the chosen mass assumption [17].
$M_{A^{\prime}}$ $M_{ST}$ 1250 1500 1750 1800 1970 2000 3000 4000 5000 6000 7000
200 4.3×$\times 10^{-4}$ 86.5×$\times 10^{-4}$ 158.0×$\times 10^{-4}$ 170.0×$\times 10^{-4}$ 180.0×$\times 10^{-4}$ 181.0×$\times 10^{-4}$ 89.5×$\times 10^{-4}$ 26.0×$\times 10^{-4}$ 6.1×$\times 10^{-4}$ 1.2×$\times 10^{-4}$ 2.3×$\times 10^{-5}$
300 9.8×$\times 10^{-7}$ 40.0×$\times 10^{-4}$ 110.0×$\times 10^{-4}$ 120.0×$\times 10^{-4}$ 142.0×$\times 10^{-4}$ 145.0×$\times 10^{-4}$ 87.4×$\times 10^{-4}$ 27.0×$\times 10^{-4}$ 6.6×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.6×$\times 10^{-5}$
400 5.4×$\times 10^{-7}$ 10.0×$\times 10^{-4}$ 65.0×$\times 10^{-4}$ 75.0×$\times 10^{-4}$ 101.0×$\times 10^{-4}$ 105.0×$\times 10^{-4}$ 79.0×$\times 10^{-4}$ 26.0×$\times 10^{-4}$ 6.7×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.8×$\times 10^{-5}$
500 3.3×$\times 10^{-7}$ 8.1×$\times 10^{-7}$ 32.0×$\times 10^{-4}$ $\times 10^{-4}$ 66.0×$\times 10^{-4}$ 69.5×$\times 10^{-4}$ 68.8×$\times 10^{-4}$ 19.0×$\times 10^{-4}$ 6.6×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.8×$\times 10^{-5}$
600 2.1×$\times 10^{-7}$ 3.5×$\times 10^{-7}$ 11.0×$\times 10^{-4}$ 17.0×$\times 10^{-4}$ 39.0×$\times 10^{-4}$ 42.0×$\times 10^{-4}$ 58.0×$\times 10^{-4}$ 23.0×$\times 10^{-4}$ 6.3×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.8×$\times 10^{-5}$
700 1.4×$\times 10^{-7}$ 1.9×$\times 10^{-7}$ 1.2×$\times 10^{-4}$ 4.1×$\times 10^{-4}$ 19.0×$\times 10^{-4}$ 22.0×$\times 10^{-4}$ 48.0×$\times 10^{-4}$ 20.0×$\times 10^{-4}$ 5.9×$\times 10^{-4}$ 1.4×$\times 10^{-4}$ 2.8×$\times 10^{-5}$
800 9.0×$\times 10^{-8}$ 1.2×$\times 10^{-7}$ 1.3×$\times 10^{-7}$ 3.2×$\times 10^{-7}$ 7.1×$\times 10^{-4}$ 9.1×$\times 10^{-4}$ 39.0×$\times 10^{-4}$ 18.0×$\times 10^{-4}$ 5.4×$\times 10^{-4}$ 1.3×$\times 10^{-4}$ 2.7×$\times 10^{-5}$
900 6.0×$\times 10^{-8}$ 7.8×$\times 10^{-8}$ 1.2×$\times 10^{-7}$ 1.4×$\times 10^{-7}$ 1.1×$\times 10^{-4}$ 2.1×$\times 10^{-4}$ 30.0×$\times 10^{-4}$ 16.0×$\times 10^{-4}$ 5.0×$\times 10^{-4}$ 1.2×$\times 10^{-4}$ 2.6×$\times 10^{-5}$
Table 2: The simplified model (based on Einstein Cartan gravity) cross-section measurements times branching ratios (in pb) calculated for different sets of the masses $M_{A^{\prime}}$ (in GeV), and $M_{ST}$ (in GeV), for the mass assumption given in table 1, with dark matter mass ($M_{\chi}=500$ GeV), the following couplings constants $\texttt{g}_{\eta}=0.2,\leavevmode\nobreak\ \texttt{g}_{D}=1.2$ and at $\sqrt{s}=13.6$ TeV.
[…]
In this model, there are several free parameters, including the mass of the torsion field ($M_{ST}$), the mass of the dark gauge boson ($M_{A^{\prime}}$), the mass of the dark matter ($M_{\chi}$), and the coupling constants ($\texttt{g}_{\eta}$ and $\texttt{g}_{D}$). As mentioned earlier, the values of these coupling constants are set at $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$. By using this combination of coupling values, we can achieve the highest possible cross-section times branching ratios. These model parameters are listed in Table 1.
We are currently investigating the potential production of a heavy neutral dark gauge boson ($A^{\prime}$) at the LHC, with a mass ($M_{A^{\prime}}$) that exceeds 100 GeV. According to the mass requirement specified in table 1, the dark matter mass must be substantial. Previous searches conducted in [50, 51] have indicated that there is a significant chance of detecting dark matter with masses between approximately 100 and 10000 GeV using various search techniques. As a result, we have designated the dark matter mass as $M_{\chi}=500$ GeV.
The typical signature of this process consists of a pair of opposite-sign muons from the decay of A plus a large missing transverse momentum due to the stable dark matter $\chi$. The events we have studied are in the following topology: $\mu^{+}\mu^{-}+E^{miss}_{T}$. This is a clean channel for Standard Model (SM) backgrounds due to the optimization of the CMS detector for this decay channel. Additionally, according to reference [33], the fake rate of the QCD background was much smaller in the dimuon channels compared to the dielectron channel.
III Simulation of signal samples and SM backgrounds
The SM background processes yielding muon pairs in the signal region are Drell-Yan (DY $\rightarrow\mu^{+}\mu^{-}$) production, the production of top quark pairs ($\text{t}\bar{\text{t}}\rightarrow\mu^{+}\mu^{-}+2b+2\nu$), $tW\rightarrow\mu^{+}\mu^{-}+2\nu+b$, and production of diboson ($W^{+}W^{-}\rightarrow\mu^{+}\mu^{-}+2\nu$, $ZZ\rightarrow\mu^{+}\mu^{-}+2\nu$ and $W^{\pm}Z\rightarrow\mu^{\pm}\mu^{+}\mu^{-}+\nu$). The second type of background is the jet background, which comes from the misidentification of jets as muons, where a jet or multijet passes the muon selection criteria. This kind of background originates from two processes: W+jets and QCD multijet. The contamination of single and multijet backgrounds in data is usually estimated from data using a so-called data-driven method, which is explained in [33]. Nevertheless, they are irrelevant to our study because our analysis is based on MC simulations only, and no events from W+jets pass the analysis pre-selection presented in the following section.
The signal samples, for the simplified model (based on Einstein-Cartan gravity), and the corresponding SM processes have been generated using MadGraph5_aMC@NLO [54] interfaced to Pythia 8 for parton shower model and hadronization [55], and DELPHES [56] for a fast detector simulation of the CMS experiment. They were generated from proton-proton collisions at the Large Hadron Collider at 13.6 TeV center of mass energy, which corresponds to the circumstances of run 3, with muon $p_{T}>10$ GeV and $|\eta|<3$ rad.
For the simplified model, with the use of mass assumption given in table 1, table 2 indicates the cross-section measurements times branching ratios calculated for different sets of the dark gauge boson (A) and torsion field ($ST$) masses. The simulated signals, used in this analysis, are private production samples, at which we used the matrix element event generator MadGraph5 aMC@NLO v2.6.7 [54].
All the Monte Carlo samples utilized in this analysis and their respective cross sections were computed in next-to-leading order and privately generated. Thus, the contributions of the signal samples and the SM background processes have been estimated from the Monte Carlo simulations, at which they are normalized to their corresponding cross-section and integrated luminosity of 52 fb-1 (the amount of certified data which is collected by CMS experiment so far during LHC run 3) [52].
IV Event selection
The selection of the event, for the analysis, has been designed to reconstruct a final state with two high transverse momentum $(p_{T})$ muons in association with missing transverse momentum, accounting for the DM candidate. The selection is made using cuts applied to different kinematic parameters. Each of the two muons should pass the following pre-selection:
$\bullet$ $p^{\mu}_{T}$ (GeV) $>30$,
$\bullet$ $\eta^{\mu}$ (rad) $<$ 2.4,
[…]
IsolationVarRhoCorr represents the isolation cut in DELPHES software to reject muons produced inside jets. In this cut, it is required that the scalar $p_{T}$ sum of all muon tracks within a cone of $\Delta R=0.5$ around the muon candidate, excluding the muon candidate itself, should not exceed 10% of the $p_{T}$ of the muon. This cut has been corrected for the pileup effect.
Thus, each event has been selected with two opposite-charge muons, and the invariant mass of the dimuon is bigger than 60 GeV since we are looking for a resonance in the high mass regime.
Figure 2: The measured dimuon invariant mass spectrum, for events passing selection 1 summarized in table 3, for the estimated SM backgrounds and different choices of dark gauge boson (A) masses generated based on the simplified model, with mass of torsion field ($M_{ST}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
Figure 3: The distribution of the missing transverse momentum, after selection 1; for the expected SM backgrounds, and different A masses produced by the simplified model, with the mass of torsion field ($M_{ST}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
Figure 2 shows the distribution of the dimuon invariant mass for events passing selection 1, summarized in table 3; The cyan histogram represents the Drell-Yan background, the yellow histogram represents the vector boson pair backgrounds (WW, WZ, and ZZ), and the red histogram represents the $t\bar{t}$ background. These histograms are stacked. While the signals of the simplified model in the framework of Einstein-Cartan gravity, which have been generated with different masses of the neutral dark gauge boson A with fixed values of the torsion field mass ($M_{ST}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV), are represented by different colored lines, and are overlaid. The corresponding distribution of the missing transverse momentum is presented in Figure 3. It is clearly shown from these figures that the signal samples are overwhelmed by the backgrounds. So, it is necessary to apply a tighter set of cuts to discriminate signals from SM backgrounds, as will be explained in the next paragraph.
In addition to selection 1, extra-tight cuts have been designed to reduce the SM background further. These tight cuts are based on three variables: the first variable is related to the invariant mass of the dimuon, at which we restricted the invariant mass of the dimuon to a small range around the mass of the dark gauge boson A to suppress backgrounds that do not peak at the A mass, a requirement that $250.9\times M_{A^{\prime}}<M_{\mu^{+}\mu^{-}}<M_{A^{\prime}}+$ 25 as suggested in [17]. The second is the relative difference between the transverse momentum of dimuon $(p_{T}^{\mu^{+}\mu^{-}})$ and the missing transverse momentum $(E^{\text{miss}}_{T})$, it has been selected to be less than 0.4. (i.e. $|p_{T}^{\mu^{+}\mu^{-}}-E^{\text{miss}}_{T}|/p_{T}^{\mu^{+}\mu^{-}}<0.4$). The third one is $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}_{T}}$, which is defined as difference in the azimuth angle between the dimuon direction and the missing transverse momentum direction (i.e. $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}_{T}}=|\phi^{\mu^{+}\mu^{-}}-\leavevmode\nobreak\ \phi^{miss}|$ ), it has been selected to be greater than 2.6 rad. These variables effectively reject DY and top quark processes. These adjustments are fine-tuned to maximize the sensitivity of the signal for a range of DM processes.
For dimuon events passing selection 1, we present the distributions of two variables for the signal presentation of the simplified model corresponding to Einstein-Cartan gravity, along with SM backgrounds. The first variable is $|p_{T}^{\mu^{+}\mu^{-}}-E_{T}^{\text{miss}}|/p_{T}^{\mu^{+}\mu^{-}}$, shown in Figure 4(a), while the second variable is $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}_{T}^{\text{miss}}}$, shown in Figure 4(b). The model signals were generated with dark gauge boson mass $M_{A^{\prime}}$ = 200 and 900 GeV, torsion field mass $M_{ST}=2000$ GeV, and dark matter mass $M_{\chi}=500$ GeV. The distributions are scaled to one. The vertical dashed lines in both figures correspond to the chosen cut value for each variable, and these tight cuts have been applied to significantly reduce the SM backgrounds.
Figure 4: Distributions of $|p_{T}^{\mu^{+}\mu^{-}}-E_{T}^{miss}|/p_{T}^{\mu^{+}\mu^{-}}$ (a) and $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}_{T}^{\text{miss}}}$ (b) for two signals presentation of the simplified model corresponding to the Einstein-Cartan gravity with $M_{A^{\prime}}=$ 200 and 900 GeV and SM backgrounds, for dimuon events with each muon is passing the pre-selection cuts listed in table 3. The vertical dashed lines correspond to the chosen cut value for each variable. All histograms are normalized to unity to highlight qualitative features.
Selection 1 Selection 2 Selection 3
$p^{\mu}_{T}>$ 30 GeV $p^{\mu}_{T}>$ 30 GeV $p^{\mu}_{T}>$ 30 GeV
$\eta^{\mu}<$ 2.4 rad $\eta^{\mu}<$ 2.4 rad $\eta^{\mu}<$ 2.4 rad
IsolationVarRhoCorr $<$ 0.1 IsolationVarRhoCorr $<$ 0.1 IsolationVarRhoCorr $<$ 0.1
$M_{\mu^{+}\mu^{-}}>$ 60 GeV $M_{\mu^{+}\mu^{-}}>$ 60 GeV $M_{\mu^{+}\mu^{-}}>$ 60 GeV
$|p_{T}^{\mu^{+}\mu^{-}}-E_{T}^{\text{miss}}|/p_{T}^{\mu^{+}\mu^{-}}<$ 0.4 $|p_{T}^{\mu^{+}\mu^{-}}-E_{T}^{\text{miss}}|/p_{T}^{\mu^{+}\mu^{-}}<$ 0.4
$\Delta\phi_{\mu^{+}\mu^{-},\vec{E}_{T}^{\text{miss}}}>$ 2.6 rad $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}_{T}^{\text{miss}}}>$ 2.6 rad
$250.9\times M_{A^{\prime}}<M_{\mu^{+}\mu^{-}}<M_{A^{\prime}}+$ 25
Pre-selection Semi-final selection Final selection
Table 3: Summary of cut-based event selections used in the analysis.
Figure 5: The dimuon invariant mass spectrum, for events passing selection 2 listed in table 3, for the estimated SM backgrounds and different choices of dark gauge boson (A) masses generated based on the simplified model, with mass of torsion field ($M_{ST}=2000$ GeV) and dark matter mass ($M_{\chi}=500$ GeV).
The dimuon invariant mass spectrum is shown in Figure 5 for events that meet the criteria listed in Table 3 (selection 2). The histograms represent the estimated SM backgrounds and various dark gauge boson (A) masses generated based on the simplified model, with a torsion field mass of $M_{ST}=2000$ GeV and dark matter mass of $M_{\chi}=500$ GeV.
V Results
The shape-based analysis has been used based on the missing transverse momentum distributions ($E^{\text{miss}}_{T}$), which are good discriminant variables since the signal distributions are characterized by relatively large $E^{\text{miss}}_{T}$ values compared to SM backgrounds. After applying selection 3 listed in table 3, the distribution of the missing transverse momentum is illustrated in figure 6 for the expected SM backgrounds and one signal benchmark corresponding to the Einstein-Cartan gravity with $M_{A^{\prime}}=200$ GeV is superimposed.
Figure 6: The distribution of the missing transverse momentum, for events passing selection 3 presented in table 3, for the expected SM backgrounds and one signal benchmark corresponding to the Einstein-Cartan gravity with $M_{A^{\prime}}=200$ GeV is superimposed.
To calculate the significance of the signal, we used the Asimov formula given in [57],
$S=\sqrt{2\times\Big{(}(N_{s}+N_{b})log\Big{(}1+\frac{N_{s}}{N_{b}}\Big{)}-N_{s}\Big{)}},$ (3)
$N_{s}$ and $N_{b}$ represent the number of signals and the total number of SM background events that pass the selections (3) mentioned in tables 3. In figure 7, the significance of A discovery potential is plotted versus the integrated luminosity for the simplified model based on Einstein-Cartan gravity, with the dark matter mass fixed at $M_{\chi}=500$ GeV, $M_{ST}=2000$ GeV and the coupling constants $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$. In the run3-LHC experiment, there is a possibility of detecting evidence for the torsion field with an integrated luminosity of more than 9 fb-1 for $M_{ST}=2000$ GeV and $M_{A^{\prime}}=200$. Furthermore, for $M_{ST}=2000$ GeV and $M_{A^{\prime}}=700$, a $5\sigma$ discovery can be achieved with an integrated luminosity exceeding 38 fb-1, as illustrated in figure 7. However, according to the discovery potential based on the integrated luminosity, achieving a $5\sigma$ evidence is not feasible when the $M_{A^{\prime}}$ masses exceed 700 GeV due to the significant decrease in the signal cross sections.
Figure 7: The significance was presented with the integrated luminosity for Einstein-Cartan gravity, considering different dark boson masses $(M_{A^{\prime}})$. The values of the coupling constants were fixed at $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$, with $M_{ST}=2000$ GeV, and the dark matter mass set at $M_{\chi}=500$ GeV. The red dotted line refers to the $5\sigma$ discovery potential.
Figure 8: 95% CL upper limits on the cross-section times the branching ratio (expected), as a function of the mediator’s mass ($M_{ST}$) based on Einstein-Cartan model, with the muonic decay of the A. The black line represents the Einstein-Cartan gravity with $M_{A^{\prime}}=200$ GeV.
Figure 9: The 95% CL upper limits on the product of the cross-section and branching fraction from the inclusive search, for variations of pairs of the simplified model parameters ($M_{ST}$ and $M_{A^{\prime}}$). The filled region indicates the upper limit. The dotted black curve indicates the expected exclusions for the nominal A cross-section.
To make a statistical interpretation for our results, we performed a statistical test based on the profile likelihood method, with the use of the modified frequentist construction CLs [58, 59] used in the asymptotic approximation [57] to derive exclusion limits on the product of signal cross sections and branching fraction Br($A^{\prime}$ $\rightarrow\mu\mu$) at 95% confidence level. The limits and significances presented have an adhoc flat 10% uncertainty systematic uncertainties taken into account to cover all possible systematic effects.
The 95% upper limit on the cross-section times the branching ratio versus the mass of torsion field $M_{ST}$, for the simplified model based on Einstein-Cartan gravity, is presented in figure 8, with the muonic decay of the A and coupling constant values of $\texttt{g}_{\eta}=0.2$ and $\texttt{g}_{D}=1.2$ and dark matter mass $M_{\chi}=500$ GeV. The black solid curve represents the simplified model for $M_{A^{\prime}}=200$ GeV. Based on figure 8, we exclude the torsion field $(ST)$ production in the mass range between 1369 - 6241 GeV as shown from the expected median.
For the simplified model (based on Einstein-Cartan gravity), the cross-section times the branching ratio limit is presented in Figure 9 as a function of the mediator’s masses $M_{ST}$ and the masses of the dark neutral gauge boson $M_{A^{\prime}}$. The region between the respective pair of the expected 95% dotted line is excluded. The results from the inclusive signal regions exclude expected values of up to $62411369<M_{ST}<$ 6241 GeV.
VI Summary
We have proposed a methodology to search for the dark neutral gauge bosons (A) at the LHC, which is produced in association with dark matter $(\chi)$ in the framework of $U(1)_{D}$ simplified model based on the Einstein-Cartan gravity. The presented analysis has been performed using the simulated proton-proton collisions corresponding to the LHC run 3 with 13.6 TeV center of mass energy, for an integrated luminosity of 52 fb-1, which corresponds to the amount of certified data collected by the CMS experiment so far during LHC run 3. Results from the muonic decay mode of A are discussed, with fixing the values of the coupling constants to be $g_{D}=1.2$, $g_{\eta}=0.2$, and dark matter mass ($M_{\chi}=500$ GeV). We have considered the variations of several parameters of the signal model: the torsion field mass $(M_{ST})$ and the dark neutral gauge boson mass $(M_{A^{\prime}})$.
The final analysis indicates that achieving a $5\sigma$ discovery of the A dark gauge boson in the $\mu^{+}\mu^{-}$ decay channel is feasible. This can be accomplished for $M_{A^{\prime}}$ values ranging from 200 to 700 GeV, given $M_{ST}=2000$ GeV, with an integrated luminosity of less than 40 fb-1 at the 13.6 TeV center of mass energy of the LHC run 3.
In case no new physics has been discovered, the general version of the analysis, which uses only event-level kinematic variables, excludes models with $62411369<M_{ST}<$ 6241 GeV at 95% confidence level (CL).
Acknowledgements.
The author of this paper would like to thank Cao H. Nam, the author of [17], for his useful discussions about the theoretical models, and for sharing with us the Universal FeynRules Output (UFO) for the model that was used for the generation of the events. This paper is based on works supported by the Science, Technology, and Innovation Funding Authority (STIF) under grant number 48289.
Data Availability Statement: This manuscript has no associated data or the data will not be deposited.